OpenStudy (anonymous):

pls help me with precal :( Find an equation in standard form for the hyperbola with vertices at (0, ±6) and asymptotes at y = ±three divided by four times x.

OpenStudy (anonymous):

we can do this

OpenStudy (anonymous):

ready ?

OpenStudy (anonymous):

actually this one is not that bad do you know what this sucker looks like ?

OpenStudy (anonymous):

uhhhh what do you mean what it looks like? what a hyperbola with a vertice at 0,+/-6 and asymptotes at the given point look like? yea

OpenStudy (anonymous):

ok so you know that the center is \((0,0)\) and the form of it must be \[\frac{y^2}{a^2}-\frac{x^2}{b^2}=1\] the \(y^2\) term comes first because of the they way it is oriented 

OpenStudy (anonymous):

the fact that the vertices are at \((0,6)\) and \((0,-6)\) tells you it looks like \[\frac{y^2}{6^2}-\frac{x^2}{b^2}=1\] and all you need is \(b\)

OpenStudy (anonymous):

so how do i find b? should i just plug in numbers and test?

OpenStudy (anonymous):

you have another piece of information asymptotes at \(y=\pm\frac{3}{4}x\)

OpenStudy (anonymous):

typo there i meant since \(a=6\) and \(\frac{a}{b}=\frac{3}{4}\) you get \(b=8\)

OpenStudy (anonymous):

ohhhh ok

OpenStudy (anonymous):

and that is done now

OpenStudy (anonymous):

Ok thanks man :^)

OpenStudy (anonymous):

yw